Cross 1: wildtype x white eyes/yellow bodies Fi: Females- 80 wildtype Males- 121 white eyes/yellow bodies The Fı generation is selfed producing: F2: wildtype White eyes/yellow bodies Wildtype eyes/yellow bodies White eyes/wildtype bodies Females 164 101 Males 132 109 3 How can I produce a chi-square test to test if the genes are sex linked and non-linked?
Answer #1

Here,

H0 : Here gens are not sex - linked.

Ha : Here genes are pro - sex - linked.

Here observed Table is

Females Males Total
WT 164 132 296
WE/YB 101 109 210
WE/YB 2 3 5
WTE/WTB 8 5 13
Total 275 249 524

Now expected taable is

Expected value for each cell is calculated by i.e.

Wildtype female Expected value = Total females * total wildtype only /total number of people = 275 * 296/524 = 155.34

WIldtype Male Expected value = Total males * total wildtype only /total number of people = 249 * 296/524 = 140.66

SO observed table is

Females Males Total
WT 155.34 140.66 296
WE/YB 110.21 99.79 210
WE/YB 2.62 2.38 5
WTE/WTB 6.82 6.18 13
Total 275 249 524

Now Value of Chi-square X2   = Σ(Oi - Ei)2/Ei

Females Males Total
WT 0.48 0.53 1.02
WE/YB 0.77 0.85 1.62
WE/YB 0.15 0.16 0.31
WTE/WTB 0.20 0.22 0.43
Total 1.60 1.77 3.37

Here X2 = 3.37

Degree of freedom = (2 -1) * (4 - 1) = 3

alpha = 0.05

X2critical = 7.8147

p - value = 0.3374 > 0.05

So we fail to reject the null hypothesis and can say that the genes are not sex linked.